OpenStudy (anonymous):

Find the area of a rhombus with a side of length 9.4 m and a vertex angle that measures 45 degrees. Round your answer to the nearest hundredth.

OpenStudy (dumbcow):

 Area = 9.4*h = (9.4^2)/sqrt2